OpenStudy (animegirl1):

If x=8, y=20. Find y when x=42

OpenStudy (anonymous):

Well, 42 divided by 8 is 5.25 and 5.25 times 20 is 105.

OpenStudy (anonymous):

I think that's correct.

OpenStudy (anonymous):

So, x=42 and y=105

OpenStudy (texaschic101):

set it up as a proportion 8/20 = 42/x now solve for x. Can you do that ?

OpenStudy (anonymous):

Or that.

OpenStudy (anonymous):

8y=42*20 8y=840 y=105

OpenStudy (texaschic101):

correct :)